# DualNILM: Energy Injection Identification Enabled Disaggregation with Deep Multi-Task Learning

### v1

@@ Abstract
Non-Intrusive Load Monitoring (NILM) offers a cost-effective method to obtain fine-grained appliance-level energy consumption in smart homes and building applications. However, the increasing adoption of behind-the-meter energy sources, such as solar panels and battery storage, poses new challenges for conventional NILM methods that rely solely on at-the-meter data. The injected energy from the behind-the-meter sources can obscure the power signatures of individual appliances, leading to a significant decline in NILM performance. To address this challenge, we present DualNILM, a deep multi-task learning framework designed for the dual tasks of appliance state recognition and injected energy identification in NILM. By integrating sequence-to-point and sequence-to-sequence strategies within a Transformer-based architecture, DualNILM can effectively capture multi-scale temporal dependencies in the aggregate power consumption patterns, allowing for accurate appliance state recognition and energy injection identification. We conduct validation of DualNILM using both self-collected and synthesized open NILM datasets that include both appliance-level energy consumption and energy injection. Extensive experimental results demonstrate that DualNILM maintains an excellent performance for the dual tasks in NILM, much outperforming conventional methods.
Index Terms:
Non-Intrusive Load Monitoring, behind-the-meter energy injection, deep multi-task learning.I Introduction
@@
Non-Intrusive Load Monitoring (NILM) technique enables disaggregation of total household energy consumption into individual appliance usage patterns [1], offering a cost-effective solution for fine-grained energy monitoring. This capability is increasingly critical for smart grid applications, enabling targeted energy-saving recommendations, supporting demand response programs, and enhancing overall grid efficiency [2]. Traditional NILM methods leverage unique electrical signatures of appliances to infer their operating states from aggregate measurements at the utility meter [3, 4, 5]. The value of accurate appliance-level monitoring becomes particularly pronounced during peak demand periods, where granular consumption insights enable utilities to implement precise demand response strategies and maintain grid stability without costly infrastructure upgrades [6].
@@
However, the rapid adoption of behind-the-meter renewable energy systems, particularly residential solar photovoltaic (PV) installations, fundamentally challenges the core assumptions of conventional NILM methods. Modern residential PV systems’ energy injection patterns can completely dominate household consumption profiles during peak generation periods [7, 5]. This paradigm shift invalidates the traditional assumption that aggregate power measurements represent purely consumptive loads, as behind-the-meter generation creates bidirectional power flows that obscure appliance signatures [8].
@@
The resulting “signal eclipse” phenomenon manifests through multiple interconnected challenges. At the signal level, routine appliance operations become indistinguishable from generation fluctuations. For instance, a refrigerator’s compressor activation, normally visible as a distinct step increase in aggregate power, vanishes when offset by simultaneous solar generation. At the temporal level, natural solar variability driven by cloud transients, atmospheric conditions, and diurnal patterns introduces non-stationary variations that mimic appliance switching events. This creates a perfect storm of ambiguity where traditional pattern recognition catastrophically fails. At the system level, when generation exceeds consumption, the net power flow reversal transforms the mathematical structure of the disaggregation problem from a constrained optimization with unique solutions to an under-determined inverse problem with infinite feasible solutions.
@@
The technical complexity of this challenge cannot be overstated. When PV generation exceeds instantaneous consumption, the disaggregation problem undergoes a fundamental mathematical transformation. Traditional NILM benefits from non-negativity constraints that bound the solution space. Each appliance contributes positively to the aggregate, creating a convex optimization landscape. Behind-the-meter generation shatters these constraints, expanding the feasible region to include arbitrary combinations of positive consumption and negative generation that yield identical net measurements. This transformation is further complicated by the electrical characteristics of modern grid-tied inverters, which operate at near-unity power factors to maximize generation efficiency. While optimal for energy production, this operational mode eliminates the reactive power signatures that traditionally help distinguish different load types, removing a critical disambiguation channel just when it is most needed.
@@
Figure 1 empirically demonstrates these theoretical challenges through controlled laboratory experiments. Our micro-inverter injection setup reveals how even modest energy generation fundamentally distorts the aggregate signal morphology. During injection periods (shown in green), the expected additive structure of load aggregation breaks down entirely. The aggregate signal (orange) deviates dramatically from the consumption-only baseline (blue), while the actual appliance states (bottom panels) become completely obscured. This controlled demonstration validates that the challenge transcends simple noise addition; it represents a fundamental incompatibility between conventional NILM architectures and the bidirectional power flows increasingly common in modern residential settings.
@@
To address these challenges, we propose DualNILM, a deep multi-task learning framework specifically designed for simultaneous appliance state recognition and energy injection disaggregation. Our key insight is that generation and consumption, rather than being antagonistic signals, contain complementary information that can be leveraged for mutual disambiguation. The framework integrates sequence-to-point learning for discrete appliance state detection with sequence-to-sequence modeling for continuous injection estimation, unified through a Transformer-based architecture that captures both short-term switching events and long-term generation trends. Critically, DualNILM natively processes multi-channel feature sequences, incorporating active power, reactive power, and extensible to additional electrical signatures. This design exploits cross-dimensional correlations to resolve ambiguities intractable in single-channel analysis. The attention mechanism dynamically weights these feature channels based on the instantaneous generation-consumption balance, effectively performing context-aware feature selection.
@@
Systematic characterization of the behind-the-meter challenge: We provide rigorous analysis of how residential-scale PV injection fundamentally transforms the NILM problem from constrained optimization to under-determined inverse problem, identifying specific failure modes including signal eclipse effects, temporal aliasing, and constraint relaxation pathologies.
Novel dual-task architecture with adaptive multi-feature fusion: We present DualNILM, which jointly optimizes appliance state recognition and energy injection estimation through complementary neural pathways. The architecture’s native multi-channel processing with attention-based feature weighting enables robust disaggregation even when generation completely dominates consumption signals.
Comprehensive validation across diverse operational regimes: Through experiments spanning controlled laboratory environments (500W micro-inverter) and synthetic augmentation with realistic residential-scale generation (2kW PV), we demonstrate DualNILM’s ability to effectively identify appliances’ state and the behind-the-meter injection.
Open research infrastructure: We release the synthetic PV-augmented versions (with realistic weather data) of REDD and UK-DALE datasets and the corresponding benchmark suites to facilitate continued research in this critical area.
The significance of this work extends beyond addressing an immediate technical challenge. With global residential solar capacity projected to exceed 5.5 TW by 2030, representing a threefold increase from 2023 levels [9, 10], the ability to maintain load visibility despite bidirectional power flows becomes critical infrastructure for grid intelligence. The dual-task paradigm we propose provides a scalable framework naturally extensible to emerging energy scenarios involving battery storage, vehicle-to-grid systems, and peer-to-peer energy trading. By demonstrating that generation and consumption can be simultaneously tracked using existing smart meter infrastructure, without requiring expensive smart inverters or dedicated generation meters, this work democratizes access to advanced energy analytics and accelerates the transition toward a truly intelligent and sustainable energy future.
The rest of the paper is organized as follows. Section II discusses and gives the literature on related work. Section III presents the problem formulation and highlights the challenge. Section IV details our proposed DualNILM framework. Section V describes the experimental setup, including the realistic energy injection data simulation, and Section VI presents the experimental results and analysis. Finally, Section VII concludes the paper and discusses future directions.
@@ II Related work
II-1 Evolution of NILM Technique
@@
II-2 Categories of NILM Methods
@@
NILM methods can also be categorized based on their task-oriented objectives. Many approaches focus on event-based detection, where the key aim is to recognize appliance switching events (ON/OFF transitions) and label those events with the corresponding device [20, 21]. This category typically relies on detecting changes in aggregate power that exceed a certain threshold, followed by a classification stage to identify which appliance caused the event. Various improvements to event detection have been proposed, including hybrid schemes that combine primary event triggers with secondary verification algorithms to reduce false detections [22]. By isolating transition points in the time series, event-based NILM methods can be highly interpretable and computationally efficient, but they may struggle with appliances that exhibit gradual power variations or multiple internal states [23]. On the other hand, the steady-state-based approaches emphasize continuous monitoring of power consumption, estimating the exact load of each device over time. These NILM methods often employ more complex models (e.g., HMMs or deep neural networks) and can capture multi-level operating states of appliances [24].
II-3 NILM Methods amid RE Integration
@@
Previous work addressing NILM with renewable energy integration is limited. In [25], the authors investigate methods to manage power injection systems in renewable setups, emphasizing system stability. Similarly, Aygen et al. [26] consider a zero-sequence current injection technique for mitigating unbalanced loads in systems integrated with renewables, which, however, is not directly applicable to NILM. Only a few studies have attempted to explicitly disaggregate solar generation from the net metered signal. Noori et al. [27] recently examined solar disaggregation and highlighted that incorporating external context like solar irradiance can improve the detection of solar generation. In [8], the authors investigate solar energy disaggregation but rely on high-frequency data and additional features, which may not be practical in real-world deployments. Other studies have considered the use of additional sensors or prior knowledge of the energy injection [28], which limits their applicability.
Our work differs from existing approaches by explicitly addressing the challenge of energy injection in NILM using a deep multi-task learning framework. By simultaneously performing appliance state recognition and energy injection disaggregation, our proposed DualNILM effectively handles the complexities introduced by RE sources without relying on additional sensors or impractical data requirements.
III Challenge and Problem Formulation
This section clarifies the underlying NILM challenge when a portion of the household power flow originates from behind-the-meter sources, such as micro-inverters or PV and battery storage. We start by revisiting the traditional NILM setup and note its known complexity. Subsequently, we describe how negative (export) flows from renewable micro-inverters break the standard constraints, broadening the search space and hindering standard disaggregation techniques. Then, we introduce the DNN-based NILM with extra features and propose a multi-task formulation—covering both conventional appliances and a dedicated injection channel—that motivates modern deep-learning solutions to this harder scenario.
III-A Motivation and Overview
@@
Traditional NILM approaches aim to infer appliance-level power consumption from aggregated household measurements, typically under the assumption that all loads are strictly nonnegative [1, 13]. In practice, however, many residences now include small-scale renewable generation (e.g., rooftop solar). The micro-inverter responsible for turning this generation into usable AC power can appear as a “negative appliance” from the meter’s perspective whenever produced electricity exceeds local consumption.
In our laboratory studies, the introduction of even a modest 500 W micro-inverter often distorted the aggregate load profile enough to render certain appliance signatures indistinct (see Fig. 1). This distortion led to substantially higher disaggregation error for methods that assume strictly nonnegative loads, as evidenced in Table III. Moreover, the difficulty can be exacerbated by low sampling rates or the lack of multi-feature data, both common in real-world metering systems [14, 5].
III-B Classic Single-Blind NILM
Let be the aggregate power at time , and assume there are distinct appliances. Denote as the on/off state of the -th appliance and as its power consumption when active. In this classical viewpoint, the total load is
where is measurement noise. Equivalently, over a horizon of length ,
with and . Since NILM typically focuses on a smaller subset of target appliances, the remainder of the loads can be aggregated into a single term , as in
Solving for is known to be NP-hard in general [1], but the purely nonnegative structure has historically allowed heuristic or machine-learning-based methods to achieve acceptable performance [3, 4, 2].
III-C When Negative Flows Disrupt Classical NILM
As household renewable penetration increases, at least one “appliance” (the micro-inverter) can feed power back to the grid, producing negative net load. The aggregated signal then becomes
where the -th device’s power, , may be negative at some time points.
III-C1 Why Behind-the-Meter Injection Complicates NILM (a Theoretical Signal Processing Perspective)
Broken Uniqueness: Nonnegative factorizations often depend on convexity-based uniqueness results; negative entries undermine these results and can make standard disaggregation fall back on naive assumptions.
Consequently, traditional NILM methods that treat all loads as nonnegative typically fail to disentangle consumption from behind-the-meter generation when measurements dip below zero or remain subdued for extended intervals. Figure 1 highlights how the micro-inverter depresses the net power curve, obscuring appliance signatures; Table III further illustrates that classical NILM approaches suffer markedly when negative flows are not accounted for, especially under limited feature sets or learning capacity. More discussion of the theoretical analysis is given in Appendix A.
III-D DNN-Based Approaches and Richer Feature Spaces
Deep learning methods, ranging from different artificial neural networks, have demonstrated promise for NILM by exploiting additional data dimensions (e.g., voltage, current, reactive power) or learning more expressive latent representations [14, 15, 30]. Formally, let denote supplementary measurements at time (with being the total feature dimension if one counts active power as well). A neural model can then be structured as
where are trainable parameters. In scenarios where the meter data are more comprehensive (e.g., capturing phase angles or higher sampling rates), the negative injection component can be learned as a distinct pattern. However, when only low-dimensional features are available, the challenge described in Section III-C remains severe, necessitating specialized modeling.
III-E A Multi-Task Formulation for Injection and Appliance Disaggregation
A practical way to handle negative flows is to treat the micro-inverter (or any injecting source) as a separate “task” alongside conventional appliances. Specifically, if total devices are considered, define the -th as the injection device with . Then, a hybrid multi-task network can produce:
where explicitly models the behind-the-meter source. This architecture allows leveraging correlations between appliance load patterns and injection events (e.g., solar generation peaks), enforcing physical constraints such as capacity limits, and ultimately improving interpretability.
III-F Key Takeaways
1.
Classical Single-Blind NILM: Cast as an NP-hard but structured problem in which nonnegative loads provide some conic geometric constraints.
2.
Behind-the-Meter Injection: Relaxing nonnegativity for at least one source broadens the solution set, complicates identifiability, and necessitates a refined treatment in order to avoid large errors.
3.
Towards a Multi-Task Solution: Leveraging multi-feature data, explicit modeling of the injecting source, and modern deep architectures can partially restore identifiability and improve real-world performance.
The next section introduces our proposed methodology, which exploits these insights by combining sequence-to-point and sequence-to-sequence deep learning strategies while explicitly modeling injection as a dedicated task.
@@ IV DualNILM
To address the challenges outlined in Section III, we propose DualNILM, a novel deep multi-task learning framework for simultaneous energy disaggregation and injection identification. This section details the architecture of DualNILM, its key components, and the motivations behind our design choices. More implementation details can be found in Appendix. B.
IV-A Method Overview and Design Rationale
In Section III, we established that behind-the-meter negative power flows from micro-inverters expand the solution space beyond the typical nonnegative constraint, compounding disaggregation ambiguity. DualNILM mitigates this difficulty by:

### v2

@@ Abstract
Non-Intrusive Load Monitoring (NILM) offers a cost-effective method to obtain fine-grained appliance-level energy consumption in smart homes and building applications. However, the increasing adoption of behind-the-meter (BTM) energy sources such as solar panels and battery storage poses new challenges for conventional NILM methods that rely solely on at-the-meter data. The energy injected from the BTM sources can obscure the power signatures of individual appliances, leading to a significant decrease in NILM performance. To address this challenge, we present DualNILM, a deep multi-task learning framework designed for the dual tasks of appliance state recognition and injected energy identification. Using a Transformer-based architecture that integrates sequence-to-point and sequence-to-sequence strategies, DualNILM effectively captures multiscale temporal dependencies in the aggregate power consumption patterns, allowing for accurate appliance state recognition and energy injection identification. Extensive evaluation on self-collected and synthesized datasets demonstrates that DualNILM maintains an excellent performance for dual tasks in NILM, much outperforming conventional methods. Our work underscores the framework’s potential for robust energy disaggregation in modern energy systems with renewable penetration. Synthetic photovoltaic augmented datasets with realistic injection simulation methodology will be open-sourced after review.
I Introduction
@@
Non-Intrusive Load Monitoring (NILM) technique enables disaggregation of total household energy consumption into individual appliance usage patterns [1], offering a cost-effective solution for fine-grained energy monitoring. This capability is increasingly critical for smart grid applications, facilitating targeted energy-saving recommendations, demand response programs, and enhanced grid efficiency [2].
@@
Conventional NILM methods operate on a key assumption: the aggregate power measurements from the utility meter represent the sum of purely consumptive loads. Then these methods leverage unique electrical signatures of appliances to infer their operating states (or power) within the total load [3, 4, 5]. However, the rapid adoption of behind-the-meter (BTM) renewable energy systems, particularly solar photovoltaics (PV), fundamentally disrupts this premise. The injection of energy from BTM sources creates bidirectional power flows that can dominate the aggregate power profile (especially during peak generation periods) [6, 5], severely obscuring the signatures of individual appliances and challenging the foundation of traditional NILM.
@@
The resulting phenomenon of “signal eclipse” manifests itself through multiple interconnected challenges. At the signal level, routine appliance operations become indistinguishable from generation fluctuations. For instance, a refrigerator’s compressor activation, normally visible as a distinct step increase in aggregate power, can be entirely offset by a simultaneous dip in solar generation. At the temporal level, natural solar variability, driven by cloud transients, atmospheric conditions and diurnal patterns, introduces non-stationary variations that mimic appliance switching events. This creates a perfect storm of ambiguity where traditional pattern recognition catastrophically fails. At the system level, when generation exceeds consumption, the net power flow reverses. This changes the disaggregation problem from a geometrically constrained optimization within the non-negative orthant to a severely under-constrained inverse problem spanning the full real space.
@@
The technical complexity of this challenge is significant. When PV generation exceeds instantaneous consumption, the disaggregation problem undergoes a fundamental mathematical transformation. Traditional NILM benefits from non-negativity constraints that bound the solution space. Each appliance contributes positively to the aggregate, creating a convex optimization landscape. In contrast, BTM generation shatters these constraints, expanding the feasible region to include arbitrary combinations of positive consumption and negative generation that yield identical net measurements. This transformation is further complicated by the electrical characteristics of modern grid-tied inverters, which operate at near-unity power factors to maximize generation efficiency. While optimal for energy production, this operational mode eliminates the reactive power signatures that traditionally help distinguish different load types. This critical disambiguation channel is lost just when it is needed most.
@@
Figure 1 empirically demonstrates these theoretical challenges through controlled laboratory experiments. Our micro-inverter injection setup reveals how even modest energy generation fundamentally distorts the aggregate signal. During injection periods (shown in green), the expected additive structure of load aggregation breaks down. The aggregate signal (orange) deviates dramatically from the consumption-only baseline (blue), while the actual appliance states (bottom panels) become completely obscured. This controlled experiment confirms that the challenge transcends simple noise addition, revealing a fundamental incompatibility between conventional NILM architectures and the bidirectional power flows of modern energy systems.
@@
To address these challenges, we propose DualNILM, a deep multi-task learning framework specifically designed for the simultaneous disaggregation of appliance states and energy injection. Our key insight is that generation and consumption, rather than being antagonistic signals, contain complementary information that can be leveraged for mutual disambiguation. The framework unifies sequence-to-point learning for discrete appliance state detection with sequence-to-sequence modeling for continuous injection estimation within a Transformer-based architecture. This design captures both short-term appliance switching events and long-term generation trends. A critical feature of DualNILM is its native processing of multi-channel feature sequences, incorporating active power, reactive power and other extensible signatures, to exploit cross-dimensional correlations and resolve ambiguities intractable in single-channel analysis. Furthermore, an attention mechanism dynamically weights these feature channels based on the instantaneous generation-consumption balance, effectively performing context-aware feature selection under dynamic conditions.
@@
Systematic characterization of BTM challenges: We analyze how behind-the-meter PV injection transforms the NILM problem from a constrained optimization to an under-determined inverse problem, formally identifying key failure modes such as signal eclipse and temporal aliasing.
Dual-task architecture design: We propose DualNILM, a framework that jointly optimizes appliance recognition and energy injection estimation. Its attention-based multi-channel feature fusion mechanism dynamically leverage complementary information across electrical signatures, enabling robust disaggregation under eclipsed signals.
Comprehensive empirical validation: We demonstrate DualNILM’s state-of-the-art performance across diverse operational regimes, from controlled laboratory experiments (500W micro-inverter) to synthetic residential-scale scenarios (2kW PV systems).
Open research infrastructure: We release the synthetic PV-augmented versions (with realistic weather data) of REDD and UK-DALE datasets and the corresponding benchmark suites to support future research in this critical area.
@@ II Related work
Evolution of NILM Technique
@@
Categories of NILM Methods
@@
NILM methods can also be categorized according to their task-oriented objectives. Many approaches focus on event-based detection, where the key aim is to recognize appliance switching events (ON/OFF transitions) and label those events with the corresponding device [16, 17]. This category typically relies on detecting changes in aggregate power that exceed a certain threshold, followed by a classification stage to identify which appliance caused the event. Various improvements to event detection have been proposed, including hybrid schemes that combine primary event triggers with secondary verification algorithms to reduce false detections [18]. By isolating transition points in time series, event-based NILM methods can be highly interpretable and computationally efficient, but may struggle with appliances that exhibit gradual power variations or multiple internal states [19]. On the other hand, the steady-state-based approaches emphasize continuous monitoring of power consumption, estimating the exact load of each device over time. These NILM methods often employ more complex models (e.g., HMMs or deep neural networks) and can capture multi-level operating states of appliances [20].
NILM Methods amid RE Integration
@@
Previous work addressing NILM with the integration of renewable energy is limited. In [21], the authors investigate methods for managing power injection systems in renewable setups, emphasizing the stability of the system. Similarly, Aygen et al. [22] consider a zero-sequence current injection technique to mitigate unbalanced loads in systems integrated with renewables, which, however, is not directly applicable to NILM. Only a few studies have attempted to explicitly disaggregate solar generation from the net metered signal. Noori et al. [23] recently examined solar disaggregation and highlighted that incorporating an external context, such as solar irradiance, can improve the detection of solar generation. In [24], the authors investigate solar energy disaggregation but rely on high-frequency data and additional features, which may not be practical in real-world deployments. Other studies have considered the use of additional sensors or prior knowledge of the energy injection [25], which limits their applicability.
Our work differs from existing approaches by explicitly addressing energy injection challenges through deep multi-task learning. DualNILM jointly performs appliance state recognition and injection disaggregation without requiring additional sensors or high-frequency data. Critically, given the scarcity of NILM datasets containing renewable generation, we also provide synthetic augmentation methodology/datasets that incorporate realistic PV injection patterns derived from actual meteorological data, enabling reproducible research in this emerging domain.
III Problem Definition
@@
Traditional NILM approaches infer appliance-level power consumption from aggregated household measurements under the assumption that all loads are strictly non-negative [1, 9]. However, modern residences increasingly incorporate BTM renewable generation, where micro-inverters inject power back to the grid, fundamentally violating this core assumption.
III-A NILM with BTM Energy Injection
In classical NILM, aggregate power at time for appliances follows:
where denotes appliance state, represents power consumption when active, and is measurement noise. Since NILM typically monitors a subset of target appliances, this becomes:
where aggregates remaining loads and noise. While solving for is NP-hard [1], the non-negative structure historically enables acceptable performance through heuristic and learning-based methods [3, 4].
With the BTM injection, the aggregate signal transforms to:
where represents energy injection. This modification fundamentally alters the problem structure: the feasible solution space expands beyond the non-negative orthant, breaking geometric constraints that classical factorization methods rely upon. The resulting identifiability problem manifests itself as increased ambiguity in explaining power dips, degraded solution stability, and insufficient discriminative information in single-channel measurements. (See Appendix A for detailed theoretical analysis.)
III-B DNN-based NILM with Dual Tasks
Deep neural networks (DNNs) have shown promise in NILM by exploiting additional features or electrical signatures beyond active power [10, 11]. Let denote the multi-dimensional feature vector at time , including active power, reactive power, current harmonics, etc. A neural model processes these features as:
where represents trainable parameters. These richer feature spaces provide crucial discriminative information, as appliances and injection sources exhibit distinct signatures across multiple electrical dimensions. However, directly applying standard architectures to the injection-augmented problem remains challenging due to the different nature of consumption versus generation patterns.
Recognizing that energy injection exhibits distinct temporal and electrical characteristics from appliance consumption, we propose treating injection estimation as a dedicated task within a multi-task learning target:
in which the following dual tasks are simultaneously tackled:
Appliance State Recognition: Identify ON/OFF states for target appliances , using learned representations that capture appliance-specific switching patterns and steady-state signatures.
Energy Injection Disaggregation: Estimate injection power , through a specialized pathway that accounts for their continuous nature and reduces the solution search space to pursue a stable optimization landscape.
This formulation enables the model to learn complementary representations: discrete state transitions for appliances and continuous power variations for injection. By explicitly modeling injection rather than treating it as negative consumption, the framework naturally accommodates bidirectional power flows while reducing optimization complexity. The dual-task NILM formulation is also well-suited for DNN-based NILM models that can leverage rich features, capture complex patterns and dependencies in the data, allowing accurate estimation of the appliance states and the energy injection [10, 11, 26].
The following section presents the architecture of our approach that implements this dual-task paradigm via specialized neural pathways, each optimized for its respective disaggregation challenge.
@@ IV DualNILM
To address the challenges described, we propose DualNILM, a novel deep multi-task learning framework for simultaneous energy disaggregation and injection identification. This section details the architecture of DualNILM, its key components, and the motivations behind our design choices. More implementation details can be found in the Appendix B.
IV-A Design Rationale
The behind-the-meter negative power flows from micro-inverters expand the solution space beyond the typical nonnegative constraint, compounding disaggregation ambiguity. DualNILM follows the Dual-Task NILM manner and mitigates this difficulty by: